A service station charges a battery using a current of 5.1A for 7.0h . How much charge passes through the battery?
Answer #1

q = I * t = 5.1 * 7 * 60 * 60 = 128520 C
Answer #2
q = 5.1 * 7 * 60 * 60 = 128520 Columbs
Answer #3

q = i.t = 5.1 x 7 x 3600 = 128520 columbs